# Positive maps and extendibility hierarchies from copositive matrices

### v1

@@ 6. Maps associated to graphs
    positive indecomposable maps, when $G\neq K_{n}$, for all values of the parameter
@@
$J-t(A_{G}\oplus A_{H})=(R_{1}+R_{2})+(E_{1}+E_{2}+\underbrace{J-J\oplus J}_{\in\operatorname{\mathsf{EWP}}}),$
@@ 6.1. Triangle-free and cyclic graphs
For the reverse implication, we first consider the case of connected triangle-free graphs. Now assume that for a connected triangle-free graph $G$ on $n$ vertices, we have $\sigma(G)=2$, i.e., ${J}-2A_{G}\in\mathsf{SPN}_{n}$. We conclude using [BSM03, Theorem 2.140], stating that if $A=(A_{ij})\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ satisfies ${\rm diag}(A)=I_{n}$, $A_{ij}\geq-1$ for all $i,j$, and if $G_{-1}(A)$ is connected, then $A\in\mathsf{SPN}_{n}$ if and only if $G_{-1}(A)$ is bipartite. (Here $G_{-1}(A)$ denotes the graph with $n$ vertices such that $A_{G_{-1}(A)}(i,j)=1$ if and only if $A_{ij}=-1$). Indeed, the entries of the matrix $A=J-2A_{G}$ are all larger than or equal to 1-1 and the graph associated to the 1-1 values ${G}_{-1}(J-2A_{G})=G$ is connected. Hence, the graph $G$ is bipartite. If the graph has more than $m$ connected components $G_{i}$, by Proposition 6.7, we have $\operatorname{min}(\sigma(G_{i}))=2$. Also note that $\sigma(G_{i})\leq\omega(G_{i})/((\omega(G_{i})-1)\leq 2$. This implies that $\forall i\,\,(\sigma(G_{i}))=2$ implying that the connected graph $G_{i}$ is bipartite. Hence $G=\bigsqcup G_{i}$ is bipartite. ∎
@@
For the cyclic graph $C_{n}$, the group of automorphisms is the dihedral group of permutations, $D_{n}$ [Rod14]. The set of $n\times n$ matrices that commute with the automorphisms, $\mathsf{C}_{G}$ are real, symmetric and circulant matrices. Since the circulant matrices are parametrized by the vector $a\in\mathbb{R}^{n}$ with $a_{i}=A_{0i}$, we use the map $\operatorname{circ}(a)=A$ for any circulant matrix $A$ associated to the vector $a$. A circulant matrix has eigenvalues given by a multiple of the discrete Fourier transform of this vector [Dav79]
${\forall k\in[n]\,;\,\lambda_{k}=\sqrt{n}(\mathcal{F}a)_{k}=\sum_{l}a_{k}\exp\Big(\frac{2\pi i}{n}kl\Big.)}$
@@
Since the set $\{r\in\mathbb{R}^{n}:r_{0}=1,\,\mathcal{F}r\geq 0\}$ is convex, we may examine its extreme points, namely $\mathcal{F}^{-1}|0\rangle=\sum_{j=0}^{n-1}|j\rangle$ and $\frac{1}{2}\mathcal{F}^{-1}(|k\rangle+|-k\rangle)=\sum_{j=0}^{n-1}\cos\big(\frac{2\pi jk}{n}\big.)|j\rangle$ ($k\neq 0$), as candidates for optimal solutions to the above problem. Consequently, we can compute the value,
@@ 6.3. Strongly regular graphs
We shall now use the symmetry properties of (rank 3) SRGs to provide a simpler characterization of the parameter $\sigma$. By the proposition Proposition 6.8, we have $\sigma(G)=\hat{\sigma}(G)$ where
@@
For the complete graph $K_{n}$, we have
$\sigma(K_{n})=\omega(K_{n})/(\omega(K_{n})-1)=n/(n-1)$
@@
The complete graph $K_{n}$ is a rank $3$ SRG with parameters (n,n1,n1,n1)(n,n-1,n-1,n-1). ∎
@@
where $\sigma(G)$ is computed using Theorem 6.13. Note that the Hamming graph of parameters (2,4), which is the direct product of two complete graphs $K_{4}$, is a strongly regular graph of rank 3 having parameters $(16,6,2,2)$; these parameters are identical to those of the Shrikhande graph from Fig. 3, proving that these two SRGs having the same parameters are non-isomorphic.
@@ 7. Bosonic extendibility and SOS hierarchies
It is known that checking copositivity is a co-NP complete problem [MK87, DG14]. Hierarchies of polynomially-sized semidefinite programs (SDP) or sum-of-squares (SOS) programs can be used to approximate the cone $\mathsf{COP}_{n}$. These hierarchies provide inner and outer approximations that converge to the actual cone. See, for example, the works of [DKP02] and [SV24] for details on such approximation schemes. First of all, recall the hierarchy of approximations to $\operatorname{\mathsf{COP}}_{n}$ introduced in Section 6:
@@
Let us recall the following facts about this hierarchy:
@@
7.1. Star extendibility
We use $\bigvee^{r}\mathbb{C}^{n}$ to denote the $r$-fold symmetric subspace of $(\mathbb{C}^{n})^{\otimes r}$.
@@ 7.1. Star extendibility
We first characterize the notion of bosonic-extendibility witness $W\in(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}$, inspired by the results in [DPS04]. To this end, let us denote by $\Pi_{S}$ the projection of $(\mathbb{C}^{n})^{\otimes r}$ onto the symmetric subspace $\bigvee^{r}\mathbb{C}^{n}$.
Proposition 7.3.
Let $W\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})\otimes\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})$ be a bipartite Hermitian operator. Then $W\in(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}$ if and only if, there exist psd matrices $Z_{0},Z_{1},\ldots,Z_{r}\in\mathcal{M}_{n}(\mathbb{C})^{\otimes(r+1)}$ such that, for all vectors $v,w\in\mathbb{C}^{d}$,
@@
$=\langle vw^{\otimes r}|Z_{0}|vw^{\otimes r}\rangle+\sum_{j=1}^{r}\langle v\bar{w}^{\otimes j}w^{\otimes r-j}|Z_{j}|v\bar{w}^{\otimes j}w^{\otimes r-j}\rangle,$
@@
Proof.
($\implies$) For any state $\tilde{\rho}\in\mathsf{PSD}(\mathbb{C}^{n}\otimes\bigvee^{r}\mathbb{C}^{n})$ that is PPT across all bipartitions, we have
$\operatorname{Tr}((W\otimes I^{\otimes(r-1)})\tilde{\rho})=\operatorname{Tr}(\rho W)\geq 0$
[…]
Let $\mathcal{C}_{0}:=\mathsf{PSD}(\mathbb{C}^{n}\otimes\bigvee^{r}\mathbb{C}^{n})$ and $\mathcal{C}_{i}:=\{X\in\mathcal{M}_{d}(\mathbb{C})^{\otimes(r+1)}\mid X^{\Gamma_{[i]}}\geq 0\}$ for $i=1,\ldots,r$. Then by the previous statement, we have $\tilde{\rho}\in\bigcap^{r}_{i=0}\mathcal{C}_{i}$ and hence $(W\otimes I^{\otimes(r-1)})\in(\bigcap^{r}_{i=0}\mathcal{C}_{i})^{\circ}=\mkern 1.5mu\overline{\mkern-1.5mu\sum^{r}_{i=0}\mathcal{C}^{\circ}_{i}\mkern-1.5mu}\mkern 1.5mu$. This implies
$(W\otimes I^{\otimes(r-1)})=\tilde{Z}_{0}+\sum^{r}_{i=1}Z_{i}^{\Gamma_{[i]}},$
where $(I\otimes\Pi_{S})\tilde{Z}_{0}(I\otimes\Pi_{S})\geq 0$ and $\forall i,Z_{i}\geq 0$. (The closure can be ignored as we can show that the set of complex polynomials of the form
$h_{Z_{0},\ldots,Z_{r}}(v,w)=\sum_{j=0}^{r}\langle v\overline{w}^{\otimes j}w^{\otimes r-j}|Z_{j}|v\overline{w}^{\otimes j}w^{\otimes r-j}\rangle,\quad v,w\in\mathbb{C}^{n},$
is closed. Indeed, this set is a convex hull of the set of homogeneous polynomials $(v,w)\mapsto|\langle\varphi|v\overline{w}^{\otimes j}w^{\otimes r-j}\rangle|^{2}$ for $\varphi\in(\mathbb{C}^{n})^{\otimes r+1}$ and $j=0,1,\ldots,r$. )
Now by define $Z_{0}:=(I\otimes\Pi_{S})\tilde{Z}_{0}(I\otimes\Pi_{S})$, it is straightforward to check
$\langle vw|W|vw\rangle\,\|w\|_{2}^{2(r-1)}=\langle vw\cdots w\,|\,W\otimes I^{\otimes(r-1)}\,|\,vw\cdots w\rangle=\langle vw\cdots w\,|\,Z_{0}+\sum_{i=1}^{r}Z_{i}^{\Gamma_{[i]}}\,|\,vw\cdots w\rangle,$
proving the implication.
[…]
$\langle vw^{\otimes r}|\big(W\otimes I_{n}^{\otimes r-1}-(Z_{0}+\sum_{j=1}^{r}Z_{j}^{\Gamma_{[j]}})\big)|vw^{\otimes r}\rangle=0$ $\quad\forall\,v,w\in\mathbb{C}^{n}$
[…]
On the other hand, we know from [Chi10, Har13] that
[…]
Let $\rho$ be an extendible state with an extension $\tilde{\rho}$ with all partitions PPT. Then,
$\operatorname{Tr}(\rho W)$ $=\operatorname{Tr}\!\left[\bigl(W\otimes I^{\otimes(r-1)}\bigr)\,\tilde{\rho}\right$]
$=\operatorname{Tr}\!\left[\bigl(I\otimes\Pi_{S}\bigr)\bigl(W\otimes I^{\otimes(r-1)}\bigr)\bigl(I\otimes\Pi_{S}\bigr)\,\tilde{\rho}\right$]
$=\operatorname{Tr}\bigl(Z_{0}\,\tilde{\rho}\bigr)+\sum_{i}\operatorname{Tr}\bigl(Z_{i}^{\Gamma_{[i]}}\,\tilde{\rho}\bigr)\geq 0,$
and therefore, $W\in\mathsf{PPTBExt}_{n}^{(r)}$. ∎
Remark 7.4.
In a similar manner, we obtain that the following are equivalent:
  1. (1)
    $W\in(\mathsf{BExt}_{n}^{(r)})^{\circ}$,
  2. (2)
    $(I_{n}\otimes\Pi_{S})(W\otimes I_{n}^{\otimes(r-1)}\bigr)(I_{n}\otimes\Pi_{S})\;\;\geq\;0$,
  3. (3)
    There exists a positive semidefinite matrix $Z\in\mathcal{M}_{n}^{\otimes(r+1)}$ such that for all $v,w\in\mathbb{C}^{n}$,
    $\langle vw\,|\,W\,|\,vw\rangle\cdot\|w\|_{2}^{2(r-1)}\;=\;\langle vw\cdots w\,|\,Z\,|\,vw\cdots w\rangle,$
Note also that $(\mathsf{BExt}_{n}^{(r)})^{\circ}\subseteq(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}\subseteq\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}$.
The following proposition mimics the result in Proposition 3.5, and provides a necessary condition for the a map to witness $r$-extendibility based on the SOS heirarchy,
@@
From Proposition 7.3, there exist psd matrices $Z_{0},\ldots,Z_{r}\in\mathcal{M}_{n}(\mathbb{C})^{\otimes r+1}$ such that
@@
Recall that we constructed positive maps by lifting copositive matrices to the matrix pairs of the form $(N,A-\mathring{N})$ in Theorem 4.1. We leave open the question whether a similar “lifting” construction of $r$-extendibility witnesses can be carried out from the hierarchies of copositive matrices.
Question 7.6.
@@
For $A\in\mathsf{K}_{n}^{(r-1)}$, do we have $X^{\mathsf{(C)LDUI}}_{{N},{A-\mathring{N}}}\in(\mathsf{PPTExt}_{n}^{(r)})^{\circ}$ for some choice of $N\in\operatorname{\mathsf{EWP}}_{n}^{\rm sa}$? Moreover is it possible to choose $\mathring{N}=\max(A)\mathring{J}$ similar to the construction in Theorem 4.1?
@@
Although the answer to the above question remains unclear, we show in the next section that another type of hierarchy of extensions, called complete graph extendibility, can be fully characterized in terms of the copositive sum of squares hierarchy.

### v4

@@ 7. Maps associated to graphs
    positive indecomposable maps, when $G\neq\mathcal{K}_{n}$, for all values of the parameter
@@
$J-t(A_{G}\oplus A_{H})=(R_{1}\oplus R_{2})+(E_{1}\oplus E_{2}+\underbrace{J-J\oplus J}_{\in\operatorname{\mathsf{EWP}}}),$
@@ 7.1. Triangle-free and cyclic graphs
For the reverse implication, we first consider the case of connected triangle-free graphs. Now assume that for a connected triangle-free graph $G$ on $n$ vertices, we have $\sigma(G)=2$, i.e., ${J}-2A_{G}\in\mathsf{SPN}_{n}$. We conclude using [SMB21, Theorem 2.140], stating that if $A=(A_{ij})\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{R})$ satisfies ${\rm diag}(A)=I_{n}$, $A_{ij}\geq-1$ for all $i,j$, and if $G_{-1}(A)$ is connected, then $A\in\mathsf{SPN}_{n}$ if and only if $G_{-1}(A)$ is bipartite. (Here $G_{-1}(A)$ denotes the graph with $n$ vertices such that $A_{G_{-1}(A)}(i,j)=1$ if and only if $A_{ij}=-1$). Indeed, the entries of the matrix $A=J-2A_{G}$ are all larger than or equal to 1-1 and the graph associated to the 1-1 values ${G}_{-1}(J-2A_{G})=G$ is connected. Hence, the graph $G$ is bipartite. If the graph has more than $m$ connected components $G_{i}$, by Proposition 7.7, we have $\operatorname{min}(\sigma(G_{i}))=2$. Also note that $\sigma(G_{i})\leq\omega(G_{i})/((\omega(G_{i})-1)\leq 2$. This implies that $\forall i\,\,(\sigma(G_{i}))=2$ implying that the connected graph $G_{i}$ is bipartite. Hence $G=\bigsqcup G_{i}$ is bipartite. ∎
@@
For the cyclic graph $C_{n}$, the group of automorphisms is the dihedral group of permutations, $D_{n}$ [Rod14]. The set of $n\times n$ matrices that commute with the automorphisms, $\mathsf{C}_{G}$ are real, symmetric and circulant matrices. Since the circulant matrices are parametrized by the vector $a\in\mathbb{R}^{n}$ with $a_{i}=A_{0i}$, we use the map $\operatorname{circ}(a)=A$ for any circulant matrix $A$ associated to the vector $a$ (we refer to Eq. 19 for the definition of circulant matrices). A circulant matrix has eigenvalues given by a multiple of the discrete Fourier transform of this vector [Dav79]
${\forall k\in[n]\,;\,\lambda_{k}=\sqrt{n}(\mathcal{F}a)_{k}=\sum_{l}a_{l}\exp\Big(\frac{2\pi i}{n}kl\Big.)}$
@@
Since the set $\{r\in\mathbb{R}^{n}:r_{0}=1,\,\mathcal{F}r\geq 0\}$ is convex, we may examine its extreme points, namely $\mathcal{F}^{-1}|0\rangle=\sqrt{n}\sum_{j=0}^{n-1}|j\rangle$ and $\frac{1}{2}\mathcal{F}^{-1}(|k\rangle+|-k\rangle)=\sqrt{n}\sum_{j=0}^{n-1}\cos\big(\frac{2\pi jk}{n}\big.)|j\rangle$ ($k\neq 0$), as candidates for optimal solutions to the above problem. Consequently, we can compute the value
@@ 7.3. Strongly regular graphs
We shall now use the symmetry properties of (rank 3) SRGs to provide a simpler characterization of the parameter $\sigma$. By Proposition 7.8, we have $\sigma(G)=\hat{\sigma}(G)$ where
@@
For the complete graph $\mathcal{K}_{n}$, we have
$\sigma(\mathcal{K}_{n})=\omega(\mathcal{K}_{n})/(\omega(\mathcal{K}_{n})-1)=n/(n-1).$
@@
The complete graph $\mathcal{K}_{n}$ is a rank $2$ SRG with parameters $(n,n-1,n-2,\mu)$, with an undefined parameter $\mu$. The result above still applies, with the positive eigenvalue $r=\lambda_{\max}(J-I)=n-1$. ∎
@@
where $\sigma(G)$ is computed using Theorem 7.13. Note that the Hamming graph of parameters (2,4), which is the direct product of two complete graphs $\mathcal{K}_{4}$, is a strongly regular graph of rank 3 having parameters $(16,6,2,2)$; these parameters are identical to those of the Shrikhande graph from Fig. 3, proving that these two SRGs having the same parameters are non-isomorphic.
@@ 8. Bosonic extendibility and SOS hierarchies
It is known that checking copositivity is a co-NP complete problem [MK87, DG14]. Hierarchies of polynomially-sized semidefinite programs (SDP) or sum-of-squares (SOS) programs can be used to approximate the cone $\mathsf{COP}_{n}$. These hierarchies either provide inner and outer approximations that converge to the actual cone. See, for example, the works of [DKP02] and [SV24] for details on such approximation schemes. First of all, we recall the hierarchy of inner approximations to $\operatorname{\mathsf{COP}}_{n}$ based on sum of squares [Par00]:
@@
The following results are known about this hierarchy:
@@
Let us mention that connection between hierarchies of semidefinite programs for polynomial optimization and hierarchies of separability tests in quantum information has been previously considered in [DPS04, FF21], where the well-known Doherty-Parrilo-Spedalieri (DPS) hierarchy for detecting entanglement has been shown to be dual to a sum-of-squares (SOS) representations for certain Hermitian polynomials. This duality has been used to establish quantitative bounds on the convergence rate of the DPS hierarchy. In this section, we establish hierarchies for LDUI/CLDUI witnesses that detect both PPT-bosonic extendibility [DPS04, NOP09] and complete graph extendibility [ACG+23], and we provide a complete characterization of such witnesses using the copositive hierarchy. In particular, we provide surprisingly wide family of entanglement witnesses which are not certifiable by any level of the DPS hierarchy, thus providing counterexamples to the question dating back to [DPS04]. We also conjecture (see Question 8.7) that the copositive hierarchy can be lifted to provide explicit witnesses for the DPS extendibility hierarchy interpolating the “lifting” in Theorem 4.1 and Theorem 5.6.
8.1. DPS hierarchy
We use $\bigvee^{r}\mathbb{C}^{n}$ to denote the $r$-fold symmetric subspace of $(\mathbb{C}^{n})^{\otimes r}$. We start by defining the DPS hierarchy, or the hierarchy of bosonic (or bose-symmetric) extendibility
@@ 8.1. DPS hierarchy
Let us first recall the following remarkable characterization of bosonic extendibility witnesses in terms of the SOS hierarchy for a certain Hermitian form. The precise connection was implicit in [DPS04, Doh14] and later made explicit in [FF21].
Theorem 8.3.
Let $W\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})\otimes\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})$ be a bipartite Hermitian operator. Then for $r\geq 1$,
  1. (1)
    $W\in(\mathsf{BExt}_{n}^{(r)})^{\circ}$ if and only if there exists a positive semidefinite matrix $Z\in\mathcal{M}_{n}(\mathbb{C})^{\otimes(r+1)}$ such that, for all vectors $v,w\in\mathbb{C}^{n}$,
    $\langle vw\,|\,W\,|\,vw\rangle\cdot\|w\|_{2}^{2(r-1)}\;=\;\langle vw\cdots w\,|\,Z\,|\,vw\cdots w\rangle.$ (32)
  2. (2)
    $W\in(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}$ if and only if there exist psd matrices $Z_{0},Z_{1},\ldots,Z_{r}\in\mathcal{M}_{n}(\mathbb{C})^{\otimes(r+1)}$ such that, for all vectors $v,w\in\mathbb{C}^{n}$,
@@
    $=\langle vw^{\otimes r}|Z_{0}|vw^{\otimes r}\rangle+\sum_{j=1}^{r}\langle v\bar{w}^{\otimes j}w^{\otimes r-j}|Z_{j}|v\bar{w}^{\otimes j}w^{\otimes r-j}\rangle,$ (33)
@@
Note that for all $r\geq 0$, we have the following inclusions:
$(\mathsf{BExt}_{n}^{(r)})^{\circ}\subseteq(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}\subseteq\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}=\{W\in\mathcal{M}^{\mathrm{sa}}_{n}(\mathbb{C})^{\otimes 2}:\langle vw|W|vw\rangle\geq 0\;\forall v,w\in\mathbb{C}^{n}\},$
where the last set is the cone of block-positive matrices. We also remark that Eqs. 32 and 33 indeed provide SOS decompositions of the Hermitian form $(v,w)\mapsto\langle vw|W|vw\rangle\|w\|_{2}^{2(r-1)}$ since they are conic combinations of Hermitian forms of the type
$v,w\in\mathbb{C}^{n}\mapsto\big|\langle\psi|v\bar{w}^{\otimes j}w^{\otimes r-j}\rangle\big|^{2},\quad\psi\in(\mathbb{C}^{n})^{\otimes r+1},\;\;j=0,1,\ldots,r.$
Now following the previous arguments in Propositions 3.5 and 5.5, we further provide a necessary condition for the LDUI/CLDUI operators to witness $r$-extendibility based on the copositive SOS hierarchy $(\mathsf{K}_{n}^{(r)})_{r\geq 0}$.
@@
From Theorem 8.3, there exist PSD matrices $Z_{0},\ldots,Z_{r}\in\mathcal{M}_{n}(\mathbb{C})^{\otimes r+1}$ such that
@@
Later in Theorem 8.11, we will establish the precise correspondence between $X^{\mathsf{LDUI}}_{{A},{B}}$ satisfying $A+A^{\top}+2{\rm Re}(\mathring{B})\in\mathsf{K}_{n}^{(r-1)}$ and the notion of complete graph extendibility.
On the other hand, it was questioned in [DPS04, Section VI] whether the set $\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}$ can be saturated by the cones $(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}$, i.e., whether the equality
$\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}=\bigcup_{r\geq 1}(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}$ (34)
holds. Since $\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}=\big(\bigcap_{r\geq 1}\mathsf{PPTBExt_{n}^{(r)}}\big)^{\circ}=\overline{\bigcup_{r\geq 1}(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}}$, the question mathematically asks whether the closure is essential or not. It was shown in [DPS04, Theorem 2] that the interior of $\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}$ is contained in $\bigcup_{r\geq 1}(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}$. In other words, every entanglement witness $W$ which is “strictly” block-positive,
$\langle vw|W|vw\rangle>0,\quad\forall\,v,w\neq 0,$
[…]
$Tr(W\,|v\rangle\langle v|^{\otimes 2})=\langle v^{\otimes 2}|X^{\mathsf{LDUI}}_{{J_{4}},{\widetilde{H}-\mathring{J}_{4}}}|v^{\otimes 2}\rangle=\langle v^{\odot 2}|\widetilde{H}|v^{\odot 2}\rangle=0$
for $v=\frac{1}{\sqrt{2}}(0,1,1,0)^{\top}\in\mathbb{C}^{4}$, and hence $W\notin{\rm int}\big(\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}\big)$.
Now we provide the following explicit ($(n^{2}-n)/2$-dimensional) counterexamples to the question in Eq. 34 which strengthens Corollary 5.9.
Theorem 8.5.
Suppose $n\geq 6$ and let $M$ be a copositive matrix which does not have $r$-SOS property for any $r$, i.e., $M\in\operatorname{\mathsf{COP}}_{n}\setminus\bigcup_{r\geq 0}\mathsf{K}_{n}^{(r)}$. Then for any $N\in\operatorname{\mathsf{EWP}}_{n}^{\mathrm{sa}}$ with $\operatorname{diag}(N)=\operatorname{diag}(M)$ and $N_{ij}\geq\frac{1}{2}M_{ij}$ $\forall\,i\neq j\in[n]$, we have
$X^{\mathsf{(C)LDUI}}_{{N},{M-\mathring{N}}}\in\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}\setminus\bigcup_{r\geq 1}(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}.$
Furthermore, all these entanglement witnesses are on the boundary of $\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}$.
Proof.
The fact $X^{\mathsf{(C)LDUI}}_{{N},{M-\mathring{N}}}\in\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}$ follows from Lemma 2.7, Proposition 3.4, and Theorem 4.1. Furthermore, Proposition 8.4 implies that $X^{\mathsf{(C)LDUI}}_{{N},{M-\mathring{N}}}\notin(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}$ for any $r\geq 0$. The last assertion follows a priori from [DPS04, Theorem 2], as mentioned earlier. ∎
Remark 8.6.
The existence of copositive matrices in $\operatorname{\mathsf{COP}}_{n}\setminus\bigcup_{r\geq 0}\mathsf{K}_{n}^{(r)}$ for arbitrary $n\geq 6$ was first shown in [LV23]. One such example is
$M=H\oplus 0_{(n-5)\times(n-5)}\in\operatorname{\mathsf{COP}}_{n}\setminus\bigcup_{r\geq 0}\mathsf{K}_{n}^{(r)},\quad n\geq 6,$
where $H\in\operatorname{\mathsf{COP}}_{5}$ is the Horn matrix defined in Eq. 3. For other examples, we refer to [LV23, Theorem 3, Example 1 and 3]. In particular, this implies that
$X^{\mathsf{(C)LDUI}}_{{J_{n}},{M-\mathring{J}_{n}}}$ $\in\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}\setminus\bigcup_{r\geq 1}(\mathsf{PPTBExt}_{n}^{(r)})^{\circ},$
$X^{\mathsf{(C)LDUI}}_{{J_{5}\oplus 0_{(n-5)\times(n-5)}},{(H-\mathring{J}_{5})\oplus 0_{(n-5)\times(n-5)}}}$ $\in\mathsf{Sep}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})^{\circ}\setminus\bigcup_{r\geq 1}(\mathsf{PPTBExt}_{n}^{(r)})^{\circ}.$
Note that the latter matrix is a (trivial) lifting of $5\otimes 5$ Hermitian matrix $X^{\mathsf{(C)LDUI}}_{{J_{5}},{H-\mathring{J_{5}}}}$ into $n\otimes n$ system (we also refer to [SN21, Proposition 4.8]). Since
$\langle vw|X^{\mathsf{LDUI}}_{{J_{5}\oplus 0_{(n-5)\times(n-5)}},{(H-\mathring{J}_{5})\oplus 0_{(n-5)\times(n-5)}}}|vw\rangle\|w\|_{2}^{2(r-1)}$
$=\Big(\sum_{1\leq i<j\leq 5}|v_{i}\overline{w}_{j}-v_{j}\overline{w}_{i}|^{2}+\sum_{i,j=1}^{5}(v_{i}\overline{v}_{j}w_{i}\overline{w}_{j}H_{ij})\Big)\Big(\sum_{j=1}^{n}|w_{j}|^{2}\Big)^{r-1}$
is SOS for no $n\geq 6$ and $r\geq 1$, this actually shows that
$\mathsf{Sep}(\mathbb{C}^{n_{A}}\otimes\mathbb{C}^{n_{B}})^{\circ}\supsetneq\bigcup_{r\geq 1}(\mathsf{PPTBExt}_{n_{A},n_{B}}^{(r)})^{\circ}$
whenever $n_{A}\geq 5$ and $n_{B}\geq 6$, by embedding $X^{\mathsf{LDUI}}_{{J_{5}},{H-\mathring{J_{5}}}}$ into proper spaces.
@@
Recall that we constructed positive maps by lifting copositive matrices to the matrix pairs of the form $(N,M-\mathring{N})$ in Theorem 4.1. We leave open the question whether a similar “lifting” construction of $r$-extendibility witnesses can be carried out from the hierarchies of copositive matrices.
Question 8.7.
@@
For $r\geq 2$ and $M\in\mathsf{K}_{n}^{(r-1)}$, do we have $X^{\mathsf{(C)LDUI}}_{{N},{M-\mathring{N}}}\in(\mathsf{PPTExt}_{n}^{(r)})^{\circ}$ for some choice of $N\in\operatorname{\mathsf{EWP}}_{n}^{\rm sa}$? Moreover, is it possible to choose $\mathring{N}=\max(M)\mathring{J}$ similar to the construction in Theorem 4.1?
@@
Note that the above assertion for $r=1$ already follows from Theorem 5.6. Although the complete answer to the above question remains unclear, we show in the next section that another type of hierarchy of extensions, called complete graph extendibility, can be fully characterized in terms of the copositive sum of squares hierarchy, generalizing some results from the literature.